OpenStudy (anonymous):

What are the roots of this quadratic equation? 2x^2 + 5x = x^2 + 7x + 35 {-5, 7} {5, -7} {2, -3} {-2, 3}

OpenStudy (luffingsails):

Try simplifying the equation into the standard form: ax^2 + bx + c = 0 Then go from there.

OpenStudy (luffingsails):

Once you do that, you'll find it much easier to work with.

OpenStudy (amistre64):

or trial and error, test the options to see what fits ...

OpenStudy (anonymous):

5x + 2x + 7x = 7x + 7x =14x @GracieFerrell

OpenStudy (luffingsails):

Sorry @chasity.ambs , this is a quadratic equation. That's incorrect.

OpenStudy (anonymous):

OKAY @luffingsails just trying to help

OpenStudy (luffingsails):

Not being mean, just trying to be clear. No offense meant.